OpenStudy (anonymous):

Patsy found a top on the clearance rack marked 60% off the original price. If the original price is $48, how much will Patsy save? $6.00 $8.00 $19.20 $28.80

OpenStudy (anonymous):

60% off of original price = the price she paid after concession then, original price - the price she paid = concession (the amount she saved) So, if the amount she saved = x, and original price = $48 (given in question). "Off of" means"original price - 60% of original price" then, 60% off of 48 = x so, 48 - 60% of 48 = x => 48 - 60/100 * 48 = x => 48 - 0.6*48 = x => 48 - 28.8 = x => x = 19.2

OpenStudy (anonymous):

$19.2